### v2

# Local Intrinsic Dimensionality of Ground Motion Data for Early Detection of Complex Catastrophic Slope Failure

Abstract
Local Intrinsic Dimensionality (LID) has shown strong potential for identifying anomalies and outliers in high-dimensional data across a wide range of real-world applications, including landslide failure detection in granular media. Early and accurate identification of failure zones in landslide-prone areas is crucial for effective geohazard mitigation. While existing approaches typically rely on surface displacement data analyzed through statistical or machine learning techniques, they often fall short in capturing both the spatial correlations and temporal dynamics that are inherently present in such data. To address this gap, we focus on ground-monitored landslides and introduce a novel approach that jointly incorporates spatial and temporal information—enabling the detection of complex landslides, including multiple successive failures occurring in distinct areas of the same slope. To be specific, our method builds upon a previously developed LID-based technique that quantifies the outlyingness of a sample relative to its kinematic neighbors, known as $s$-LID. We extend its capabilities in three key ways. (1) Kinematic enhancement: we incorporate velocity into the $s$-LID computation to better capture short-term temporal dependencies and deformation rate relationships. (2) Spatial fusion: we apply Bayesian estimation to aggregate $s$-LID values across spatial neighborhoods, effectively embedding spatial correlations into the LID scores. (3) Temporal modeling: we introduce a temporal variant, $t$-LID, that learns long-term dynamics from time series data, providing a robust temporal representation of displacement behavior. Finally, we integrate both spatial and temporal LID components into a unified framework, referred to as spatiotemporal LID ($st$-LID), to identify samples that are anomalous in either or both dimensions. Extensive experiments show that $st$-LID consistently outperforms existing methods in failure detection precision and lead-time, offering valuable support for landslide early warning systems and targeted risk intervention to enhance community resilience and preparedness strategies.
1 Introduction
Local Intrinsic Dimensionality (LID) has emerged as a powerful measure for identifying anomalous or outlying behavior in high-dimensional datasets. By effectively capturing representative dimensions within data, LID has demonstrated significant successes across various scenarios such as intrusion detection for IoT networks [7], identifying adversarial attacks on neural networks [15], and outlier detection for high-dimensional data [1].
Recent advances in granular media mechanics have also demonstrated the effectiveness of LID in the early detection of impending failures in small scale laboratory tests on granular materials [26, 23, 14]. At larger scales, such failures manifest as some of the most devastating natural hazards. Among these, catastrophic landslides are becoming increasingly frequent and severe due to climate change, seismic activity, and human intervention [20, 18, 8]. They can occur suddenly with little warning, often resulting in significant infrastructure damage, loss of life, and long-term economic disruption [5, 13, 20]. This sudden onset underscores the importance of timely and accurate identification of high-risk areas to enable early intervention, save lives, and reduce property loss.
Existing works commonly apply traditional machine learning and statistical techniques, such as $K$-means clustering [21] and Empirical Dynamic Quantiles (EDQs) [19, 22], to monitoring data on surface displacement in the precursory failure regime, which has been proved to be one of the most direct indicators for identifying the location and timing of slope collapses [11, 2]. However, the high dimensionality and inherent spatiotemporal characteristics of such data often challenge the traditional detection methods which are built upon uni-variate time series analysis. As a result, the original informative, multidimensional data is frequently reduced to single dimensional data, leading to significant information loss and potentially inaccurate detection, in both spatial and temporal aspects. Specifically, clustering approaches primarily focus on the spatial detection of high-risk regions without considering the temporal dependencies inherent in the landslide processes. In contrast, studies on EDQs leverage the temporal dynamics of the displacement data but overlook the spatial correlations inside the data. Therefore, these existing works (1) fail to fully exploit both spatial and temporal dependencies inherited in the spatiotemporal displacement data, (2) fail to identify the sudden changes or accelerations in displacement over time. Subsequently, these limitations can lead to inaccurate detection and delay warnings, missing opportunities for timely intervention.
To address the challenge in learning with spatiotemporal data, we extend earlier studies of LID in [26] to establish a LID-based framework for dynamic detection of high-risk areas most prone to slope failure at large field scales, by accounting for the outlyingness of kinematic behavior across multiple spatial and time scales. We first add the velocity ($v_{t}$) at each time stamp into target sample, allowing the algorithm to leverage the variability of the displacement and learn the short-term temporal dependence. Incorporating velocity is crucial, as many landslide early warning frameworks rely on velocity thresholds to trigger alerts and guide response actions [4, 24]. Unlike displacement, which tracks cumulative movement, velocity reflects the rate of change—making it a more immediate indicator of accelerating instability.
Figure 1: Detection of location and order of successive failures in distinct sites of a slope. Top: The first failure event $C1$ (blue frame) occurred at $t=3385$. Bottom: Four methods detect impending failure locations at an earlier time step $t=3337$ (2 hours before collapse $C1$). Each monitoring point is colored based on the likelihood of failure at its location from 0 to 1. K-means produce binary results, 0 or 1. EDQs highlight points representative of the dynamics against the lower bound 0 for background. $s$-LID correctly identifies both areas of potential collapse but cannot distinguish their order. Only the proposed $st$-LID can detect and isolate the location of the first failure event $C1$ that is relevant at this time step.
Secondly, we propose $t$-LID to measure the outlyingness of a target sample with respect to its $t$ nearest temporal neighbors (historical records), enabling it to capture the long-term dynamics from the entire time series data. Finally, we integrate two LID-based scores to highlight the points that are outlying with respect to both spatial and temporal neighborhoods. Specifically, we apply a sigmoid function to normalize both scores into probabilistic values. Then, we integrate two probabilistic scores via multiplication to emphasize samples where both scores are high. This integration not only reduces false positives but also enables earlier detection of hazardous shifts, providing actionable insights well before visible damage occurs. We name the final score spatiotemporal LID ($st$-LID). Experimental results demonstrate that the proposed $st$-LID can pinpoint the actual landslide location with accuracy and efficiency (up to $22\times$ improvement in precision, up to $8\times$ improvement in lead-time). This highlights the practical value of $st$-LID in real-world landslide early warning systems, with the potential to save lives and protect critical infrastructures.
Our contributions can be summarized as:
  • We propose a novel latent geometry-based method called $st$-LID for spatiotemporal outlier detection. This approach integrates an extended version of existing LID method with newly proposed spatial fusion and temporal variant, capturing both spatial correlations and temporal dynamics with efficacy and efficiency.
  • We validate the proposed $st$-LID through extensive experiments on real-world datasets, demonstrating its competitive performance with respect to precision and lead-time of failure detection, and the computation efficiency of the algorithm.
  • We propose a systematic failure detection procedure based on $st$-LID, showcasing its effectiveness and efficiency in early prediction of precise failure locations. This highlights the practicability of the $st$-LID in real-world landslide early warning systems, enabling timely interventions and enhanced risk management.
2 Background & Preliminaries
Although catastrophic landslides are driven by multiple factors, surface displacement remains a reliable and measurable indicator of impending failure [11, 16, 2, 6]. In the lead-up to collapse, ground motion exhibits a distinct spatiotemporal clustering dynamics, with high-risk zones increasingly differentiating themselves from more stable areas of the slope [21, 25]. Figure 2 shows a typical example of the displacement time series corresponding to a catastrophic slope collapse. Notably, the displacement of points in the failure region manifest three stages of landslide creep (initial, steady-state and accelerated). These patterns highlight the potential for using machine learning and data mining techniques to enable the automatic early detection of slope failure based on surface displacement data (e.g., [16, 22, 2]).
Figure 2: Example of a real-world slope Failure (Top) and the displacement time series of each monitored point (Bottom). The failure (a collapse advised by domain expert) is located inside the blue frame. Each time series at the bottom represents the displacements of a monitored point along with the time. Two monitored points (inside and outside the failure region) are marked via $\times$’ with different colors in the site map (their corresponding time series are highlighted with the same colors). The red dashed line indicates the time of failure (ToF). And The purple dashed lines indicates three stages of landslide creep.
2.1 Statistical Solutions
Traditional approaches employ statistical methods like $K$-means (with $K=2$) to divide the monitored area into stable and high-risk regions based on their displacement values. Then, metrics like silhouette score and normalized mutual information (NMI) are applied to measure the clustering quality and consistency, subsequently, confirming the detection of the failure regions [21]. While $K$-means can generally highlight the high-risk regions based on the clustering outputs, its results are inherently binary and depend heavily on the class distinguishability to produce clear and stable boundaries for failure region. Additionally, without leveraging the temporal dynamics, the clustering results are solely based on displacement values at specific time step. These lead to noisy and discrete detection of failure regions, making it hard to highlight finer-scale regions of potential failure.
To account for temporal dynamics and locate specific regions, Empirical Dynamic Quantiles (EDQs) is applied to select the representative displacement time series ${\bm{x}}_{1:T}^{*(q)}$ based on the quantile level $q$:
${\bm{x}}_{1:T}^{*(q)}=\arg\min_{\{{\bm{x}}_{t}\}\in\mathbb{C}_{m}}\sum_{t=1}^{T}\rho_{q}({\bm{y}}_{t}-{\bm{1}}x_{t}).$
Where $\mathbb{C}_{m}=\{{\bm{x}}_{t}|1\leq t\leq T,{\bm{x}}_{t}\in\mathbb{R}^{m}\}$ is the candidate set, $\bm{1}$} denotes an $m$-dimensional unit vector, ${\bm{y}}_{t}$ is the observation which candidate $x_{t}$ is tested against to, and $\rho_{q}(a)=q\cdot a^{+}+(1-q)\cdot a^{-}$ measures the asymmetrical deviation with $a^{+}=\max(a,0),a^{-}=\max(-a,0)$ [19].
When selecting the top quantile levels (e.g., $q\geq 0.75$) EDQs will identify time series that are consistently reflect high percentile behaviors. Consequently, the potential failure regions can be pinpointed by these time series and their corresponding monitored points [22]. However, since EDQs only focuses on displacement at each monitored points independently without leveraging the spatial relationship between them, it can be easily biased by inherent noise of the displacement data, leading to sparsely distributed results in the monitored site which are infeasible for further analysis.
2.2 Local Intrinsic Dimensionality
While the traditional approaches directly analyzing the raw data values, recent studies that exploit the latent properties of data has proven to be more effective and efficient for machine learning and data science research [15, 3, 10]. The Intrinsic Dimensionality (ID) is one of the latent characteristics of data. It describes the minimal number of latent factors for data representation without losing significant information.
Generally, given a dataset $\mathcal{D}\in\mathbb{R}^{n\times m}$ that consists $n$ samples with $m$ dimensions. If the samples in $\mathcal{D}$ can be effectively approximated by $d$ latent dimensions, where $d<<m$, we say $\mathcal{D}$ has intrinsic dimension of $d$. Mathematically, the ID $d$ can be derived from the expansion model that measures the relationship between volume and radius in an expanding ball at the point of interest in Euclidean space:
$\frac{V_{2}}{V_{1}}=\left(\frac{r_{2}}{r_{1}}\right)^{d}\iff d=\frac{\ln(V_{2}/V_{1})}{\ln(r_{2}/r_{1})}.$
Where $V_{1,2}$ are two volumes and $r_{1,2}$ are two radius.
Similarly, when we focus on a specific sample inside a dataset, a local version of ID is measured, and the corresponding measurement is named as Local Intrinsic Dimensionality (LID). Improving upon the original expansion model, the LID model represents the volume using probability measure, allowing multi modality data sources.
Let $\mathcal{F}(r)$ be a real-valued, non-zero function around an open interval of $r\in\mathbb{R}$. If $\mathcal{F}$ is continuously differentiable at $r$, then the LID of $\mathcal{F}$ at $r$ can be defined [9]:
$\text{LID}_{\mathcal{F}(r)}:=\lim_{\Delta r\rightarrow 0^{+}}\frac{\ln(\mathcal{F}(r+\Delta r))}{\ln(1+\frac{\Delta r}{r})}=r\frac{\mathcal{F}^{\prime}(r)}{\mathcal{F}(r)},$ (1)
wherever the limit exists.
Consequently, the LID of a given queried sample $x$ can be derived from (1), letting $r>0$ be a continuous random variable representing the distance from $x$ to its neighbors (other data samples), and $\mathcal{F}(r)$ be the cumulative distribution of $r$ [26]:
$\text{LID}(x)=\lim_{r\rightarrow 0}\text{LID}_{\mathcal{F}(r)}.$ (2)
Several estimators are proposed to approximate the unknown function $\mathcal{F}$, making LID computable. Among these estimators, Maximum Likelihood Estimator (MLE) has been broadly applied due to its simplicity and efficiency. This also leads to the development of $s$-LID for failure detection in granular mechanics [26]: given a data sample at monitored point $p$ with the displacement value of $x_{t}^{p}$, MLE estimates the LID of ${\bm{x}}_{t}=\langle x_{t}^{p}\rangle$ at a fixed time step $t$ based on its distance to its $s$ nearest neighbors in kinematic space (absolute difference of displacement values):
$s\text{-LID}({\bm{x}}_{t})=-\left(\frac{1}{s}\sum_{i=1}^{s}\frac{dist_{i}({\bm{x}}_{t})}{dist_{s}({\bm{x}}_{t})}\right)^{-1},$ (3)
where $s\geq 2$ and $dist_{i}({\bm{x}}_{t})$ is the Euclidean distance of ${\bm{x}}_{t}$ to its $i^{th}$ nearest neighbor with displacement value of $x_{t}^{i}$: $dist_{i}({\bm{x}}_{t})=|x_{t}^{p}-x_{t}^{i}|$
Intuitively, $s$-LID measures the expansion rate of the neighborhood around a data sample, offering insight to its local density level. A high $s$-LID value suggests rapid distance growth with the increasing radius, reflecting a larger separation between queried sample and its nearest neighbors. This behavior corresponds to a sparser region around the queried sample, indicating that the queried sample is more outlying comparing to its neighbors. Thus, $s$-LID is applied to detect the potential failure regions (outlier) around a monitored area. Comparing to traditional methods, firstly, as a real-valued function, $s$-LID is able to produce more details than binary results from K-means by delivering different levels of outlyingness. Secondly, $s$-LID can also provide similar results to EDQ when picking up monitored locations with top $s$-LID values. However, similar to K-means, $s$-LID also focus on spatial relationship at a fixed time step, failing to leverage the temporal dynamics which has been proven to be critical in landslides analysis. Additionally, although $s$-LID can provide more detailed results for the potential failures, there are still multiply high-risk regions sparsely distributed around the monitored area, increasing the time complexity of identifying the real failure regions.
3 Methodology
To address the limitation of existing $s$-LID model on learning with spatiotemporal data and improve its practicability in landslide application, we propose a novel spatiotemporal measurement to leverage both spatial correlation and temporal dynamics inherent in the displacement data.
3.1 Leveraging Velocity in LID Calculation
Firstly, to introduce short-term temporal dependency, we extend the data sample into a vector consists both displacement value and velocity $v_{t}^{p}=x_{t}^{p}-x_{t-1}^{p}$ at monitored point $p$ and time step $t$, ${\bm{x}}_{t}=\langle x_{t}^{p},v_{t}^{p}\rangle$. And the distance of a queried sample ${\bm{x}}_{t}$ to its $i^{th}$ nearest neighbor at monitored point $i$ becomes:
$dist_{i}({\bm{x}}_{t})=\sqrt{(x_{t}^{p}-x_{t}^{i})^{2}+(v_{t}^{p}-v_{t}^{i})^{2}}.$ (4)
Applying this new distance calculation to equation (3), we get the $s\text{-LID}({\bm{x}}_{t})$ that incorporates both displacement value and velocity at given time step. This approach can effectively mitigate the bias introduced by samples exhibiting high displacement values but remain relatively stable. Figure 3 depicts the underlying idea with
Figure 3: Example high-risk areas and corresponding time series with & without velocity in $s$-LID computation. Top: Site map at last time step ($t=4983$), with two monitored points highlighted by top $s$-LID values, representing the high-risk locations. $\star$ is the $s$-LID without velocity, $\star$ is the $s$-LID with velocity. Bottom: Displacement time series for each monitored point, with the same color.
an real-world example: In the site map at top, different monitored points are selected as high-risk candidates ($\star$ and $\star$) according to the largest $s$-LID values at the last time point ($t=4983$) when queried samples with or without the velocity. The corresponding time series at bottom illustrates the different temporal dynamics between these two monitored points, reflecting the aforementioned bias when using only displacement values for $s$-LID computation. Note that the actual failure region is around the monitored point $\star$. This fact justifies our enhancement of incorporating the velocity component.
3.2 Spatial Fusion of LID
Although $s$-LID measures the query point’s characteristic against its spatial neighbors, it finds the nearest neighbors in the kinematic space, without involving the spatial dependencies in physical space. To leverage the spatial relationships of all monitored points during the LID calculation, inspired by Bayesian LID algorithm [12], we integrate the $s$-LID results in a small neighborhood of the query point, enhancing the LID score with actual spatial dynamics. Figure 4 depicts the underlying idea, where the query point (red) is updated based on its neighbors (green). Note that the actual number of neighbors is decided during the experiments.
Figure 4: Spatial fusion of $s$-LID. Red dot in the middle is the query point and its neighbors are colored with green. The final LID value of the query point is calculated based on its own observation and the pre-calculated $s$-LID of these neighbors at last time step.
Given a region with pre-calculated $s$-LID values at time step t1t-1. For each query point ${\bm{x}}_{t}$, let $\{f^{i}_{t-1}(l)\}_{i=1}^{k}$ be the $k$ neighbors’ prior densities that are built from last time step $s$-LID values. With a set of weights ${\bm{w}}=\{w_{1},\cdots,w_{k}\}$ summing up to $1$, we can combine these priors with logarithmic pooling:
$f^{pool}_{t}(l|{\bm{x}}_{t}):=\left(\int_{0}^{\infty}\prod_{i=1}^{k}(f^{i}_{t-1}(l))^{w_{i}}dl\right)\prod_{i=1}^{k}(f^{i}_{t-1}(l))^{w_{i}}.$ (5)
To simplify the calculation, we can assume Gamma densities $f^{i}=Ga(\alpha_{i},\beta_{i}),i=1,\cdots,k$, and equation (5) becomes a Gamma density as well:
$f^{pool}=Ga(\alpha_{p}=\sum_{i=1}^{k}w_{i}\alpha_{i},\beta_{p}=\sum_{i=1}^{k}w_{i}\beta_{i}).$ (6)
Here, we apply Gaussian kernel to create weights that account for the distance from query point to its neighbors:
$w_{i}=\frac{\exp\left(-\frac{dist_{i}({\bm{x}}_{t})^{2}}{2\sigma^{2}}\right)}{\sum_{j\in k}\exp\left(-\frac{dist_{j}({\bm{x}}_{t})^{2}}{2\sigma^{2}}\right)}.$
Where $dist_{i}({\bm{x}}_{t})$ is the Euclidean distance between query point and its $i^{th}$ neighbor. This weight ensures closer neighbors have higher influences of the pooling results. Then, the parameters of the Gamma density can be derived from the weighted mean and variance of the neighbors’ $s$-LID values:
$\mu$ $=\sum_{i=1}^{k}w_{i}s\text{-LID}({\bm{x}}_{t-1}^{i})$
$\sigma^{2}$ $=\sum_{i=1}^{k}w_{i}\left(s\text{-LID}({\bm{x}}_{t-1}^{i})-\mu\right)^{2}$
$\alpha_{p}$ $=\frac{\mu^{2}}{\sigma^{2}},\quad\beta_{p}=\frac{\mu}{\sigma^{2}}.$ (7)
Where $s$-LID(${\bm{x}}_{t-1}^{i}$) is the $s$-LID value of the $i^{th}$ neighbor of query point at time step t1t-1.
Once we get the current observation of query point, we can estimate its likelihood contribution via the density ratio within the $k$-neighborhood:
$\alpha_{o}=k,\beta_{o}=-\sum_{i=1}^{k}\frac{dist_{i}({\bm{x}}_{t})}{dist_{k}({\bm{x}}_{t})}.$ (8)
Finally, applying Bayesian estimation, we can get the posterior estimation of the query point’s $s$-LID with parameters from equations (7) and (8) following Gamma conjugacy:
$s\text{-LID}^{*}=\frac{\alpha_{p}+\alpha_{o}}{\beta_{p}+\beta_{o}}.$
This approach can effectively leverage spatial correlation in geophysical monitoring data to produce more robust LID estimates compared to purely local methods on kinematic space, ensuring geometric coherency and spatial continuity in LID calculation.
3.3 Temporally Informed LID
Thirdly, to learn the long-term dynamics of the time series data, we propose a temporal version of LID namely $t$-LID. Same as $s$-LID, $t$-LID is also estimated based on the distance expansion rate. Rather than spatial neighborhood, $t$-LID focuses on temporal neighborhood of the target sample, on another word, the historical values of the given sample until the analyzing time step $t$. Notice that, the raw displacement values are highly non-stationary across the temporal dimension (see Figure 2, 3). To mitigate the outlyingness contributed by such non-stationarity, we compute $t$-LID solely on the velocity values. This method also aligns with the common practice for de-trending non-stationary time series [17]. Hence, given a time series of samples at a fixed monitored point $l$: ${\bm{x}}_{1:t}=\{{\bm{x}}_{1},{\bm{x}}_{2},\cdots,{\bm{x}}_{t}\},{\bm{x}}_{t}=\langle v_{t}^{l}\rangle$, we can write $t$-LID of the queried sample at current time step ${\bm{x}}_{t}$ as:
$t\text{-LID}({\bm{x}}_{t})=-\left(\frac{1}{t}\sum_{i=1}^{t}\frac{dist_{i}({\bm{x}}_{t})}{dist_{t}({\bm{x}}_{t})}\right)^{-1},$ (9)
where $t\geq 2$ and $dist_{i}({\bm{x}}_{t})=|v_{t}^{l}-v_{i}^{l}|$ is the distance of the queried sample ${\bm{x}}_{t}$ at current time step $t$ to its $i^{th}$ nearest neighbor at a historical time step $i$.
Similar to $s$-LID, which evaluates the outlyingness of a queried sample based on its spatial neighbors, $t$-LID measures the outlyingness of the queried samples against its own past records, enabling effective outlier detection within time series data. This temporal perspective allows $t$-LID to capture the dynamic changes along with the time, identifying the significant deviations that may relate to the instability and emerging risks in a landslide process.
3.4 Spatiotemporal LID
Until now, we get two LID measurements: $s$-LID that quantifies the spatial outlyingness, and $t$-LID that captures the temporal deviations. At the final step, we propose a novel approach that integrates them into a unified spatiotemporal outlier detection framework. This integration aims to leverage the strengths of both methods on spatial and temporal outlier detection, addressing the limitations of traditional methods that consider these aspects in isolation.
To achieve an appropriate integration, we begin with the scale unification: the two LID measurements operate on different targets and has different scales, therefore, they are not directly comparable. We apply a sigmoid function $\sigma(x)=\frac{1}{1+e^{-x}}$ to effectively convert raw LID values into probabilistic scores between $[0,1]$, allowing for consistent interpretation of the outlier likelihood. Then, to ensure only the points exhibiting significant outlyingness in both spatial and temporal perspective are highlighted, a multiplicative fusion is applied. Hence, the final probabilistic score, which we name as spatiotemporal LID ($st$-LID), can be written as:
$st\text{-LID}({\bm{x}}_{t})=\sigma(s\text{-LID}({\bm{x}}_{t}))\times\sigma(t\text{-LID}({\bm{x}}_{t})).$ (10)
The proposed $st$-LID effectively leverages both spatial correlation and temporal dependence of the spatiotemporal displacement data. In contrast to the existing methods that fail to provide precise and localized details of the failure regions, $st$-LID can pinpoint the potential failure regions at early time step, enabling further analysis and timely rescue for catastrophic landslides.
3.5 Real-World Failure Detection via $st$-LID
The proposed $st$-LID is designed to identify the outliers in displacements data with precision and efficiency. However, it remains an open question to apply it in real-world scenario. In this section, we propose a systematic failure detection procedure to enable the practical deployment of $st$-LID.
Assuming a continuous monitoring system based on $st$-LID scores at each monitoring time step, we detect a potential failure region around the monitored point ${\bm{x}}^{*}\in\mathbb{R}^{2}$ at time step $t$ if for $n$ consecutive time steps $\{t-n+1,t-n+2\cdots,t\}$ the following conditions stand:
[…]
Where $MP$ is the set of all monitored points, and $\mathcal{B}_{\epsilon}({\bm{x}}_{t}^{*})=\{{\bm{y}}\in\mathbb{R}^{2}\mid\|{\bm{y}}-{\bm{x}}_{t}^{*}\|<\epsilon\}\text{, }\epsilon>0\in\mathbb{R}$ is a small open ball around ${\bm{x}}_{t}^{*}$ to allow certain fluctuation caused by noise.
This detection standard follows the thresholding of the failure ($\geq 0.5$) and ensure the convergence of the detection (small open ball). The choice of $n$ can be decided by the domain experts, it acts as a trade off between the sensitivity and the fidelity.
4 Experiments
In this section, we evaluate the proposed $st$-LID on three real-world catastrophic landslides cases, against the aforementioned existing methods on early detection of the actual failure regions.
4.1 Datasets
The datasets consist surface displacement measurements data from three real-world operating mine sites which experienced fatal collapses (failures). Due to confidential policy, we name these datasets as $M_{1}$, $M_{2}$, and $M_{3}$ respectively, without revealing their actual locations and names.
  • $M_{1}$: Contains 2622 monitored points with surface displacement measurements recorded in 8000 time steps, and there are two ground truth collapses happened at two different locations at $t=3385,7320$ respectively, we refer to them as $C1$ and $C2$.
  • $M_{2}$: Contains 5844 monitored points with surface displacement measurements recorded in 6000 time steps, and the single ground truth collapse happened at $t=5264$.
  • $M_{3}$: Contains 6624 monitored points with surface displacement measurements recorded in 6500 time steps, and the single ground truth collapse happened at $t=5806$.
The actual time interval between two consecutive time steps is $2.5$ minutes in $M_{1}$, and $6$ minutes in $M_{2}$ and $M_{3}$.
More details of these datasets including the visualization of the displacement time series are provided in the Appendix.
Figure 5: Visualization of the ground truth in 4 datasets (columns) and corresponding detection results from 5 methods (rows). Detection results are colored based on the likelihood of the failure (see color-bar for detailed range and order)
Table 1: Precision Score $Prec.$ ($\uparrow$) & Lead-time $\Delta t$ ($\uparrow$) results from all methods. $\uparrow$ represents higher values are better. Best performance are highlighted in bold format.
Metrics Methods $M_{1}$ ($C1$) $M_{1}$ ($C2$) $M_{2}$ $M_{3}$
$Prec.$ K-means 0.044 0.940 0.128 0.999
DBSCAN 0.047 0.557 0.298 0.705
LOF 0.043 0.800 0.494 0.732
EDQs N.A. N.A. N.A. N.A.
s-LID 0.045 0.945 0.514 0.845
st-LID (ours) 1.000 1.000 1.000 1.000
$\Delta t$ K-means 0 220 (9.2 hrs) 0 427 (1.8 days)
DBSCAN 0 1351 (2.4 days) 214 (21.4 hrs) 1096 (4.6 days)
LOF 0 355 (14.8 hrs) 0 127 (12.7 hrs)
EDQs 0 0 0 0
s-LID 0 1489 (2.6 days) 1383 (5.8 days) 1142 (4.8 days)
st-LID (ours) 80 (3.3 hrs) 1726 (3.0 days) 1760 (7.3 days) 1214 (5.1 days)
4.2 Experimental Setup
The target of this work is to pinpoint the real failure locations at early time step, therefore, we evaluate the methods using two metrics:
  • Precision ($Prec.$) measures the proportion of correct detection results (failure /collapse actually happened at the detected locations) relative to all detections from the method:
    $Prec.=\frac{\text{Correct Detections}}{\text{All Failure Detections}}.$
    Higher precision score reflects less false alarms, indicating the method is trustworthy.
  • Lead-time ($\Delta t$) measures the difference between the detection time of model and the actual time of failure, quantifying how early the method detects the failure regions:
    $\Delta t=t_{failure}-t_{detection}.$
    A longer lead-time represents earlier detection of failure, enabling timely invention and reactions.
We compare the proposed $st$-LID with existing failure detection methods including K-means, EDQs, and $s$-LID (extended version with velocity). We also include one cluster method: Density-based spatial clustering of applications with noise (DBSCAN), and one outlier detection algorithm: Local Outlier Factor (LOF). Due to the differences in output types across these methods, we unify the detection threshold a fair comparison. Specifically, in K-means, DBSCAN, and LOF, all monitored points from high-risk class are regarded as failure detections. For EDQs, the monitored points whose time series are selected as quantile level $p\geq 0.5$ by the algorithm are regarded as failure detections. For $s$-LID, the results are linearly converted to scores between $[0,1]$, and the failure detections are monitored points with $s\text{-LID}\geq 0.5$.
The precision score $Prec.$ are computed using the detection results at the actual time of failures (collapses). Noted here, EDQs only produces limited number of points rather than a region, therefore, it is not feasible for precision calculation. Regarding the lead-time $\Delta t$ computation, to avoid the impact of noisy results in baselines, we use the top $10$ results from each method. To be specific, these are the monitored points with: (1) $10$ shortest distance to outlying cluster’s centroid in K-means and DBSCAN, (2) $10$ highest LOF scores, (3) first $10$ quantile levels in EDQs algorithm, and (4) $10$ highest $s$-LID with scores larger than $0.5$. The proposed $st$-LID is designed to pinpoint the failure regions, thus, we use all its detection results without pruning. We record the first time when the selected points consistently fall within the actual failure regions (until the time of failure), and use it to compute $\Delta t$. We assign $\Delta t=0$ for those methods cannot pinpoint the actual failure ahead of time.
Note that we do not include the commonly used metrics Recall because it requires a reliable count of all actual failures which is typically unavailable for real-world landslide scenarios. Moreover, the objective of this work is to develop manageable and high-confident early warnings for landslide failures, rather than to maximize the number of detected events at any false alarm burden. Thus, we focus on the Precision and Lead-time only.
4.3 Experimental Results
We provide both the qualitative and quantitative results in this section via the visual analysis and the aforementioned numeric metrics. The former provides intuitive insights into failure patterns, while the latter offers rigorous evaluation of performance.
Figure 5 shows the ground truth collapse locations (blue rectangle frame, advised by the domain experts) along with the detection results from three methods, we exclude EDQs because it only outputs limited number of points which does not form any patterns (refer to Figure 1). It is obvious that the proposed $st$-LID can precisely pinpoint the actual collapse locations in all four cases while other methods picking up both collapsed locations and the noisy results in most cases.
The detailed precision scores and lead-time are listed in Table 1, from which we can observe that the proposed $st$-LID consistently achieves perfect precision score, indicating its results are trustworthy and applicable for a real-world landslide warning system. Furthermore, apart from providing the precise detections, $st$-LID can also produce timely predictions of correct failure regions well before the actual time of failure in all these cases. Considering that the baseline results are denoised while full results from $st$-LID are used in this evaluation, the comparison is inherently biased in favor of the baselines. Nevertheless, the proposed $st$-LID outperforms them in terms of the overall performance, demonstrating its effectiveness and efficiency in landslide early warning tasks.
4.4 Computation Time
Efficiency is one of the important factors in evaluation of landslide warning systems. A method loses its practical values if it requires excessive computation time, especially in real-time monitoring scenarios where timely alert are crucial for risk management. In this section, we compare the computation time of all methods.
Table 2: Computation time of each failure detection methods.
Methods Computation Time
K-means 0.192 seconds
DBSCAN 0.162 seconds
LOF 0.101 seconds
EDQs 19.66 hours
s-LID 0.649 seconds
st-LID (sequential) 5.342 seconds
st-LID (parallel) 0.721 seconds
The results are listed in Table 2. Here, we record the computation time of each method detecting the potential failures at time step $t=3337$, aligning with the provided examples in Figure 1. We provide two versions of $st$-LID because the $s$-LID integration and $t$-LID computation happens at each monitored point (each time series), thus, one can choose to sequentially compute it or parallel the computation. From the table, we can see except EDQs, which takes significant long time to compute, all three other methods are efficient and only take seconds to finish. This fact further justify the practicability of the proposed $st$-LID in real-time monitoring and analysis of the landslide-prone area.
4.5 Real-World Failure Detection Simulation
Until now, we have demonstrated the effectiveness and efficiency of the proposed $st$-LID conceptually via multiple experiments. In this section, we simulate a real-time monitoring using $M_{1}$ dataset and demonstrate a practical failure detection procedure on $C1$ collapse.
Given the facts that (1) $st$-LID produces perfect precision score, (2) actual time interval between records is $2.5$ minutes and the maximal computation time of $st$-LID is around $5$ seconds, it is practical to compute $st$-LID at each monitoring time step and perform analysis accordingly. These also validate the detection procedure in equation (3.5). Applying the proposed detection procedure on $M_{1}$ ($C1$), we demonstrate a simulation of the failure detection, the details are shown in Figure 6. Here, we use $n=10$ for higher fidelity considering
Figure 6: Simulation of real-world failure detection using stst-LID. The monitored point with the highest $st$-LID is marked by $\star$.
the real-world impact of a failure detection. From the figure, we can see the target ${\bm{x}}_{t}^{*}$ is fluctuating between two monitored points within a small region, following the small ball condition in equation (3.5). The target point ${\bm{x}}_{t}^{*}$ does not meet the detection threshold at initial steps ($t\in[3290,3294]$), but starts to consistently exceeds such threshold $st\text{-LID}({\bm{x}}_{t}^{*})\geq 0.5$ from $t=3295$. Finally, following to the detection procedure, the failure is detected at $t=3305$ with $n=10$. Note that, given the actual time interval of the dataset ($2.5$ mins), this detection is about $3.3$ hours ahead of the actual collapse at this region, allowing a timely intervention for this failure.
5 Conclusion
In this study, we introduce $st$-LID, a novel method designed to improve the early and accurate detection of failures in landslide-prone areas. Unlike existing approaches that typically focus on either spatial or temporal aspects of displacement data, $st$-LID effectively incorporate both the spatial correlations and temporal dynamics inherent in the displacement data. By extending the traditional LID framework with a short-term temporal feature (velocity), spatially fusing the LID values in a neighborhood to incorporate spatial correlations, and introducing a temporal variant to account for long-term dependencies, our method offers a unified approach to identifying outliers across both space and time. Comprehensive experiments demonstrate the effectiveness and efficiency of the proposed $st$-LID in terms of early and precise failure detection. These findings highlight the practical value of $st$-LID in real-world landslide early warning systems, paving the way for timely risk management and potentially saving lives.
6 Acknowledgment
We would like to thank the GroundProbe for providing the datasets and support for this study.
Appendix 0.A Displacement Time Series Data
We provide visualization of the displacement time series here. Since these datasets are multivariate, for a better visualization, similar to the example in Figure 2, we highlight two time series in-/outside the failure region, and use red dash line to indicate the actual failure time.
Figure 7: Visualization of the displacement time series.
Figure 7 shows the heatmap of three displacement time series, from which we can see most displacement values stay low all the time with several monitored points experience displacement increasing until their corresponding failure time.